In the design of a chemical plant, the following expenditures and revenues are estimated after the plant has achieved its desired production rate.

Total capital investment (TCI) : $10 000 000

Working capital investment (WCI) : $1 000 000

Annual sales : $8 000 000 /year Annual expenditures : $2 000 000 /year

Assuming straight-line depreciation over a 10-year project analysis period, determine

a) The return on the investment after taxes

b) The payback period


1
Expert's answer
2021-05-16T17:46:16-0400

(a)

Deducting Depreciation (straight line method)=initial cost-scrap value/estimated useful life

=1000000-0/10

=$100000

So Annual inflow=600000-100000

             =$500000 (after depreciation)

Less tax@ 21%=500000-21(500000)

            =$395000

Annual Inflow (after tax before depreciation)

=395000+100000

=$495000


As annual inflow are equal we fill find annuity discount factor first (ADF)

ADF=Initial investment /annual inflow (after tax before depreciation)

=1000000/495000

=2.02

Referring Present Value Annuity table and n=10

We get IRR=49% (approximately)

 

Net present value =$6140000

Present Value of inflow-present value of Outflow

As discount rate=0%, The value will remain the same.

PV of total inflow (Before Depreciation, After Tax) +Working Capital

=$595000*12+$100000

=$7240000

PV of Outflow = Capital Investment + Working Capital 

=1000000+100000

=$1100000

NPV=7240000-1100000

=$6140000

Rate of return=39.5%

Average inflow(after dep and tax)

=(800000-200000-100000)-105000(21% of 500000)

=$395000

"Rate \space Of\space Return=\frac{395000}{\frac{1000000}{100}}"

    =39.5%

Depreciation available in year 10=$100000 (at it is straight line method, so depreciation will remain same each year)

(b)

Payback Period=Net initial investment / Annual inflow( Before depreciation after tax)


Annual inflow= Annual Revenue - Annual Expenditure




Pay Back Period=1000000/495000

             =2.02 years